OpenStudy (jose_alamo32):

how do i do this?.... the lengths of the sides of a triangle are in a ratio of 3:6:8. the perimeter of the triangle is 510cm. find the length of the largest side.

OpenStudy (amistre64):

since the ratio stays the same; some multiple of them will produce the desired perimeter

OpenStudy (amistre64):

if our side lengths were a ratio if : a:b:c and they added up to some value k then we could determine that n(a+b+c) = k and solve for n. the side lengths would therefore be: an, bn, cn

OpenStudy (jose_alamo32):

so a=3 b=6 c=8. so i add them and get k? but how would get n?....can you tell me the answer? im not 100% on it

OpenStudy (amistre64):

k = the perimeter; so we are adding them all up to get the perimeter. n is the only unknown now. 3n + 6n + 8n = 510 or: n(3+6+8) = 510, solve for n we see that 8n is going to be the longest side:

OpenStudy (jose_alamo32):

o ok lol wow thats easy. thank you and have another one..... the measures of two supplementary angles are in the ratio 4:11 what is the measure of the smallest angle? that would be 4 right?

OpenStudy (amistre64):

the smallest angle would be 4n yes supplementary angles add up to 180 soo 4n + 11n = 180, its the same solution concept as the other one 15n = 180, therefore n=180/15 the angle is therefore: 4(180)/15

OpenStudy (jose_alamo32):

alright thank you :)

OpenStudy (amistre64):

good luck :)